OpenStudy (anonymous):

If the Pr[E] = , 10/11 determine the odds in favor of event E

OpenStudy (anonymous):

Wow.

OpenStudy (anonymous):

This is just a notation thing.

OpenStudy (anonymous):

So, odds in favor of \(X\) is written as \(m:n\) where\[ \Pr(X) = \frac m{m+n} \]

OpenStudy (anonymous):

\[ \frac{10}{11} = \frac{m}{m+n} \]This gives us two equations: \[ m = 10\\ m+n = 11\\ \implies n=1 \]So \[ m:n=10:1 \]

OpenStudy (anonymous):

i knew that

OpenStudy (anonymous):

What do you want?